# Thermalization and irreversibility of an isolated quantum system

### v2

@@ Abstract
The irreversibility and thermalization of many-body systems can be attributed to the erasure of spread non-equilibrium state information by local operations. This thermalization mechanism can be demonstrated by the sequence $[\hat{O}^{\dagger}\hat{O}(t)]^{N}$, where $\hat{O}$ is a local operator, $\hat{O}(t)=e^{i\hat{H}t}\hat{O}e^{-i\hat{H}t}$, $\hat{H}$ is the system Hamiltonian, and $N$ denotes the number of repetitions. We begin by preparing a non-equilibrium initial state with an inhomogeneous particle number distribution in a one-dimensional Hubbard model. As particles propagate and interact within the lattice, the system evolves into a highly entangled quantum state, where the entanglement entropy satisfies a volume law, yet the information of the initial state remains well preserved. The local operator $\hat{O}$ erases part of the information in the entangled state, altering the interference of the system wavefunction and the disentangling process during time-reversed evolution. Repeatedly applying $\hat{O}^{\dagger}\hat{O}(t)$ leads to a monotonic increase in the entanglement entropy until it saturates at a steady value. By incorporating this information erasure mechanism into the one-dimensional Hubbard model, our numerical simulations demonstrate that in a completely isolated system, a thermalization process emerges. Finally, we discuss the feasibility of implementing related quantum simulation experiments on superconducting quantum processors.
@@
Theoretically, in an isolated system, regardless of system size or interaction complexity, if the Hamiltonian $\hat{H}$ can be inverted as $-\hat{H}$, the system evolves backward to its initial state, preserving quantum information. Now, let us consider a thought experiment: Initially, gas particles are confined to a container’s corner, with each wave function forming a localized wave packet, exhibiting low position uncertainty and high momentum uncertainty. As the wave packets evolve over time and spread in space, collisions and scattering among the particles lead to a near-uniform state, where each particle becomes highly entangled with other particles and exhibits significant position uncertainty. To reverse the evolution and bring the gas particles back to their initial localized state, the system would have to disentangle the highly entangled wave functions through a series of collisions and reconstruct the original localized wave packets via constructive interference. In fact, from a quantum mechanical perspective, any process that leads to a decrease in thermodynamic entropy generally involves disentanglement and constructive interference. Experimentally, achieving such a reversal imposes extremely stringent conditions. Even small local perturbations $\hat{O}$ can disrupt the disentanglement and constructive interference process. In our thought experiment, interactions between the gas particles and the container walls could introduce slight perturbations, altering the phase along specific transmission paths and disrupting the coherence required for perfect reversal.
@@
$|\psi_{0}\rangle$ is the initial quantum state, $H$ is the unperturbed Hamiltonian, and $\delta H$ represents a small perturbation to the system. The Loschmidt Echo measures the overlap between the time-evolved states under $H$ and $H+\delta H$, capturing the reversibility of quantum dynamics and the stability of quantum information. A rapid decay of $L(t)$ indicates strong sensitivity to perturbations, which is often associated with quantum chaos and thermalization. Here, we focus on the information erasure effect of the sequence $[\hat{O}^{\dagger}\hat{O}(t)]^{N}$. For a typical macroscopic system, the operation $\hat{O}^{\dagger}\hat{O}(t)$ is expected to erase the spread information of the system’s initial non-equilibrium state, thereby preventing it from returning to non-equilibrium. Consequently, the application of the sequence $[\hat{O}^{\dagger}\hat{O}(t)]^{N}$ should lead to a monotonic increase in the system’s entanglement entropy, representing a mechanism for irreversibility and thermalization. The introduction of such perturbations may seem to violate the assumption of an isolated system. However, this issue can be resolved by incorporating the perturbation source, such as the container walls in the thought experiment, into the isolated system. Here, $\hat{O}(t)=e^{i\hat{H}t}\hat{O}e^{-i\hat{H}t}$ can be regarded as the time evolution or propagation of the operator $\hat{O}$ under the Hamiltonian $\hat{H}$. A similar sequence is the out-of-time-order correlator (OTOC), which serves as an important tool for studying the dynamical evolution of quantum information and the thermalization properties of quantum many-body systems [12, 13, 14, 15, 16].
@@
In this work, we construct a quantum many-body system in which the initial state information rapidly spreads. Numerical simulations demonstrate the fast-growing entanglement entropy dynamics starting from a non-equilibrium state. We then numerically show the information erasure and the resulting irreversible thermalization effects induced by the sequence $[\hat{O}^{\dagger}\hat{O}(t)]^{N}$ in this system. By introducing the corresponding local perturbations into the model, we construct an isolated quantum many-body system that exhibits rapid thermalization. Finally, we discuss the feasibility of conducting related research on superconducting quantum computing chips.
@@
Figure 2: The dynamical evolution of entanglement entropy and mutual information. (a) The evolution of entanglement entropy $S_{\rho}$ is presented for different subsystem sizes $A$. The blue, orange, and green lines correspond to $A=\{1,2,3\}$, $\{1,2\}$, and $\{1\}$, respectively. The black dashed line denotes the subsystem entropy predicted by the microcanonical ensemble. (b) The orange line represents the time evolution of the average mutual information between different lattice sites, defined as $\langle I(i:j)\rangle=\frac{1}{N}\sum_{i,j}I(i:j)$ , where $N=C_{6}^{2}$ is the number of site pairs. The orange shaded region indicates the standard deviation of mutual information, $\sigma_{I}=\sqrt{\frac{1}{N}\sum_{i,j}(I_{ij}-\langle I_{ij}\rangle)^{2}}$ The black dashed line denotes the mutual information predicted by the microcanonical ensemble. The inset shows mutual information fluctuations between different lattice sites $i$ and $j$ within the time range marked by the dashed box.
@@
Figure 3: Simulation of Information Erasure with $\hat{O}(t)$. In panels (a) and (b), the orange and blue dashed lines represent the entanglement entropy and mutual information evolution with and without local perturbations, respectively. The solid orange and blue lines are their envelopes, while the shaded areas indicate discrepancies between the two cases. The entanglement entropy $S(\rho_{A})$ is computed for subsystem $A=\{1,2,3\}$. The system evolves forward from $t=0$ to $250$ and backward from $t=270$ to $520$, with local operations applied only during $t=250$ to $270$. Panel (c) depicts the evolution of entanglement entropy and mutual information under the $[\hat{O}^{\dagger}\hat{O}(t)]^{N}$ sequence, where $t=50$, and $\hat{O}$ matches that in panels (a) and (b). Panel (d) shows the standard deviation of $\langle\hat{n}_{i\neq 2}\rangle$ and $\langle\hat{J}_{i}\rangle$ over all possible Fock states as a function of $N$. Panel (e) displays the average value of $\langle\hat{n}_{2,\tau}\rangle$ for all possible Fock states at lattice site 2, with different initial states, as a function of $N$. The blue (orange) curve corresponds to the initial state $|0\rangle_{2,\tau}$ ($|1\rangle_{2,\tau}$), and the shaded region represents the corresponding statistical standard deviation.
@@
Here, we focus on the reduction of quantum state information caused by $\hat{O}(t)$. Since the isolated system remains in a pure state, its quantum state information is quantified by the entanglement entropy. A higher entropy corresponds to less information. In our model, we examine the process of erasing system quantum state information during thermalization by applying the $[\hat{O}^{\dagger}\hat{O}(t)]^{N}$ sequence. As shown in Fig. 3(c), repeatedly applying this sequence leads to a monotonic increase in the system’s entanglement entropy until it reaches a stable value. More importantly, the mutual information between different lattices in the system steadily decreases, reaching a stable value. This stable value is related to the spread time $t$ of $\hat{O}(t)$. To illustrate the information erasure property of the $[\hat{O}^{\dagger}\hat{O}(t)]^{N}$ sequence, we examine the expectation values of the particle number and particle current at different lattice sites in the final state, considering all possible Fock states as the initial states of the system. The standard deviation of local observable expectation values across different initial states as a function of $N$ is shown in Fig. 3(d). Beyond a certain $N$, the expectation values of microscopic mechanical quantities for different initial states converge, indicating that after undergoing the $[\hat{O}^{\dagger}\hat{O}(t)]^{N}$ sequence, these quantum states become indistinguishable through local measurements. Lattice site 2, where the local operation $\hat{O}$ is applied, retains some dependence on the initial state. Specifically, if the initial state of lattice site 2 is $|0\rangle_{2,\tau}$ ($|1\rangle_{2,\tau}$), the expectation values in the final state for all possible Fock initial states converge, as shown by the blue (orange) line in Fig. 3(e), with the shaded region representing the standard deviation over different initial states.
@@
Figure 4: Isolated system model and entanglement entropy dynamics. (a) A perturbative particle $\phi$ is introduced into the model from Fig. 1(b). This particle moves only between two lattice sites within the dashed box and starts at lattice site 2. (b) Dynamical evolution of entanglement entropy and mutual information after introducing $\phi$. The entanglement entropy $S(\rho_{A})$ is computed for subsystem $A=\{1,2,3\}$, considering only the $\tau$ and $\upsilon$ particles. In (b), the orange line represents the average mutual information between lattice sites 1–6, with the shaded area indicating its standard deviation. (c) The model structure in (a) mapped onto a superconducting qubit chip. Polygons represent superconducting qubits, solid lines denote $XY$ coupling, and dashed lines indicate $ZZ$ coupling. Qubits marked with small balls are initialized in state $|1\rangle$.
@@
Summary and outlook.— From the perspective of quantum information, we demonstrate the mechanism of irreversibility and thermalization in the $[\hat{O}^{\dagger}\hat{O}(t)]^{N}$ sequence through local erasure of spread non-equilibrium initial state information. This mechanism is closely related to research on the decoherence[30], eigenstate thermalization hypothesis, quantum information scrambling[31, 32, 33, 34], and information erasure[35, 36, 37, 38], and further theoretical investigations are necessary. Future quantum simulations will enable unprecedented exploration of these phenomena.
@@ A.1 Quantum simulation on a superconducting quantum processor
Quantum simulations enable the study of the quantum dynamical evolution of the Hubbard model. Using the model in Fig. 4(a) as an example, we illustrate how such research can be conducted on a superconducting quantum chip. A 14-qubit chip can be constructed, as shown in Fig. 4(c), consisting of two 6-qubit rings forming inner and outer circles, with two additional qubits connected to the outer ring. Between neighboring qubits within each ring, there is a nearest-neighbor XY interaction, while between corresponding qubits in the inner and outer rings, there is a ZZ interaction. The other two qubits have only ZZ coupling with the qubits in the outer ring. The $XY$ interaction term corresponds to the hopping between neighboring lattice sites in our model, while the $ZZ$ interaction represents the repulsive or attractive interaction between different particles at the same site. Interaction parameters between qubits on the superconducting qubit chip can be tuned via couplers. Additionally, the coupling strength and sign between superconducting qubits can be controlled through these couplers, enabling quantum simulation of the $\hat{O}^{\dagger}e^{i\hat{H}t}\hat{O}e^{-i\hat{H}t}$ sequence evolution. There have been some experimental demonstrations of dynamical evolution in one-dimensional bosonic Hubbard models on superconducting qubit chips [26], as well as studies on thermalization and scrambling in bosonic Hubbard models on ladder structures [24, 23], and other experiments on two-dimensional bosonic Hubbard models [27, 28, 29].

### v3

@@ Abstract
The irreversibility and thermalization of many-body systems can be attributed to the erasure of spread non-equilibrium state information by local operations. This thermalization mechanism can be demonstrated by the sequence of $\hat{O}_{i}(t_{i})$, where $\hat{O_{i}}$ is a local operator, $\hat{O_{i}}(t_{i})=e^{i\hat{H}t_{i}}\hat{O_{i}}e^{-i\hat{H}t_{i}}$, $\hat{H}$ is the system Hamiltonian, $t_{i}$ can take positive, negative, or zero values, and the sequence is arranged according to the subscript $i$. We numerically demonstrate the information erasure of initial non-equilibrium quantum states through such sequence in a one-dimensional Hubbard model. During this process, the system’s entanglement entropy increases monotonically toward a stable value. By incorporating this information erasure mechanism into an isolated system, our numerical simulations reveal that in this completely isolated system, a thermalization process emerges.
@@
Theoretically, in an isolated system, regardless of system size or interaction complexity, if the Hamiltonian $\hat{H}$ can be inverted as $\hat{H}$, the system evolves backward to its initial state, preserving quantum information. Now, let us consider a thought experiment: Initially, gas particles are confined to a container’s corner, with each wave function forming a localized wave packet, exhibiting low position uncertainty and high momentum uncertainty. As the wave packets evolve over time and spread in space, collisions and scattering among the particles lead to a near-uniform state, where each particle becomes highly entangled with other particles and exhibits significant position uncertainty. To reverse the evolution and bring the gas particles back to their initial localized state, the system would have to disentangle the highly entangled wave functions through a series of collisions and reconstruct the original localized wave packets via constructive interference. In fact, from a quantum mechanical perspective, any process that leads to a decrease in thermodynamic entropy generally involves disentanglement and constructive interference. Experimentally, achieving such a reversal imposes extremely stringent conditions. Even small local perturbations $\hat{O}$ can disrupt the disentanglement and constructive interference process. In our thought experiment, interactions between the gas particles and the container walls could introduce slight perturbations, altering the phase along specific transmission paths and disrupting the coherence required for perfect reversal.
@@
$|\psi_{0}\rangle$ is the initial quantum state, $H$ is the unperturbed Hamiltonian, and $\delta H$ represents a small perturbation to the system. The Loschmidt Echo measures the overlap between the time-evolved states under $H$ and $H+\delta H$, capturing the reversibility of quantum dynamics and the stability of quantum information. A rapid decay of $L(t)$ indicates strong sensitivity to perturbations, which is often associated with quantum chaos and thermalization. Here, we focus on the information erasure effect of the sequence of $\hat{O_{i}}(t_{i})$. For a typical macroscopic system, the operation $\hat{O_{i}}(t_{i})$ is expected to erase the spread information of the system’s initial non-equilibrium state, thereby preventing it from returning to non-equilibrium. Consequently, the application of the sequence of $\hat{O_{i}}(t_{i})$ should lead to a monotonic increase in the system’s entanglement entropy, representing a mechanism for irreversibility and thermalization. The introduction of such perturbations may seem to violate the assumption of an isolated system. However, this issue can be resolved by incorporating the perturbation source, such as the container walls in the thought experiment, into the isolated system. Here, $\hat{O_{i}}(t_{i})=e^{i\hat{H}t_{i}}\hat{O_{i}}e^{-i\hat{H}t_{i}}$ can be regarded as the time evolution or propagation of the operator $\hat{O_{i}}$ under the Hamiltonian $\hat{H}$. A similar sequence is the out-of-time-order correlator (OTOC), which serves as an important tool for studying the dynamical evolution of quantum information and the thermalization properties of quantum many-body systems [12, 13, 14, 15, 16]. The essence of the OTOC sequence lies in the emergence of non-commutativity between initially commuting local operators after time evolution.
@@
In this work, we construct a quantum many-body system in which the initial state information rapidly spreads. Numerical simulations demonstrate the fast-growing entanglement entropy dynamics starting from a non-equilibrium state. We then numerically show the information erasure and the resulting irreversible thermalization effects induced by the sequence of $\hat{O_{i}}(t_{i})$ in this system. By introducing the corresponding local perturbations into the model, we construct an isolated quantum many-body system that exhibits rapid thermalization. Finally, we discuss the feasibility of conducting related research on superconducting quantum computing chips.
@@
Figure 2: The dynamical evolution of entanglement entropy and mutual information. (a) The evolution of entanglement entropy $S_{\rho}$ is presented for different subsystem sizes AA. The blue, orange, and green lines correspond to $A=\{1,2,3\}$, $\{1,2\}$, and $\{1\}$, respectively. The black dashed line denotes the subsystem entropy predicted by the microcanonical ensemble. (b) The orange line represents the time evolution of the average mutual information between different lattice sites, defined as $\langle I(i:j)\rangle=\frac{1}{N}\sum_{i,j}I(i:j)$ , where $N=C_{6}^{2}$ is the number of site pairs. The orange shaded region indicates the standard deviation of mutual information, $\sigma_{I}=\sqrt{\frac{1}{N}\sum_{i,j}(I_{ij}-\langle I_{ij}\rangle)^{2}}$ . The black dashed line denotes the mutual information predicted by the microcanonical ensemble. The inset shows mutual information fluctuations between different lattice sites $i$ and $j$ within the time range marked by the dashed box.
@@
Figure 3: Simulation of Information Erasure with $\hat{O}(t)$. In panels (a) and (b), the orange and blue dashed lines represent the entanglement entropy and mutual information evolution with and without local perturbations, respectively. The solid orange and blue lines are their envelopes, while the shaded areas indicate discrepancies between the two cases. The entanglement entropy $S(\rho_{A})$ is computed for subsystem $A=\{1,2,3\}$. The system evolves forward from $t=0$ to 250250 and backward from $t=270$ to $520$, with local operations applied only during $t=250$ to 270270. Panel (c) depicts the evolution of entanglement entropy and mutual information under the $\hat{O}^{\dagger}\hat{O}(t)]^{N}$ sequence, where $t=50$, and $\hat{O}$ matches that in panels (a) and (b). Panel (d) shows the standard deviation of $\langle\hat{n}_{i\neq 2}\rangle$ and $\langle\hat{J}_{i}\rangle$ over all possible Fock states as a function of $N$. Panel (e) displays the average value of $\langle\hat{n}_{2,\tau}\rangle$ for all possible Fock states at lattice site 2, with different initial states, as a function of $N$. The blue (orange) curve corresponds to the initial state $|0\rangle_{2,\tau}$ ($|1\rangle_{2,\tau}$), and the shaded region represents the corresponding statistical standard deviation.
@@
Here, we focus on the reduction of quantum state information caused by $\hat{O_{i}}(t_{i})$. Since the isolated system remains in a pure state, its quantum state information is quantified by the entanglement entropy. A higher entropy corresponds to less information. In our numerical simulation, we examine the process of erasing system quantum state information during thermalization by applying a sequence $\hat{O}^{\dagger}\hat{O}(t)]^{N}$. With fixed $\hat{O}$ and $t$ here, non-commutativity of the time evolution operators is introduced by inserting $\hat{O}^{\dagger}$. As shown in Fig. 3(c), applying this sequence leads to a monotonic increase in the system’s entanglement entropy until it reaches a stable value. More importantly, the mutual information between different lattices in the system steadily decreases, reaching a stable value. This stable value is related to the spread time $t$ of $\hat{O}(t)$. To illustrate the information erasure property of the $\hat{O}^{\dagger}\hat{O}(t)]^{N}$ sequence, we examine the expectation values of the particle number and particle current at different lattice sites in the final state, considering all possible Fock states as the initial states of the system. The standard deviation of local observable expectation values across different initial states as a function of $N$ is shown in Fig. 3(d). Beyond a certain $N$, the expectation values of microscopic mechanical quantities for different initial states converge, indicating that after undergoing the $\hat{O}^{\dagger}\hat{O}(t)]^{N}$ sequence, these quantum states become indistinguishable through local measurements. Lattice site 2, where the local operation $\hat{O}$ is applied, retains some dependence on the initial state. Specifically, if the initial state of lattice site 2 is $|0\rangle_{2,\tau}$ ($|1\rangle_{2,\tau}$), the expectation values in the final state for all possible Fock initial states converge, as shown by the blue (orange) line in Fig. 3(e), with the shaded region representing the standard deviation over different initial states. It can be inferred that in a general sequence of $\hat{O_{i}}(t_{i})$, there is no issue of initial state information retention.
@@
Figure 4: Isolated system model and entanglement entropy dynamics. (a) A perturbative particle $\phi$ is introduced into the model from Fig. 1(b). This particle moves only between two lattice sites within the dashed box and starts at lattice site 2. (b) Dynamical evolution of entanglement entropy and mutual information after introducing $\phi$. The entanglement entropy $S(\rho_{A})$ is computed for subsystem $A=\{1,2,3\}$, considering only the $\tau$ and $\upsilon$ particles. In (b), the orange line represents the average mutual information between lattice sites 1–6, with the shaded area indicating its standard deviation. (c) The model structure in (a) mapped onto a superconducting qubit chip. Polygons represent superconducting qubits, solid lines denote XYXY coupling, and dashed lines indicate ZZZZ coupling. Qubits marked with small balls are initialized in state $|1\rangle$.
@@
Summary and outlook.— From the perspective of quantum information, we demonstrate the mechanism of irreversibility and thermalization in the $\hat{O_{i}}(t_{i})$ sequence through local erasure of spread non-equilibrium initial state information. This mechanism is closely related to research on the decoherence[26], eigenstate thermalization hypothesis, quantum information scrambling[27, 28, 29, 30], and information erasure[31, 32, 33, 34], and further theoretical investigations are necessary. Future quantum simulations will enable unprecedented exploration of these phenomena.
@@ A.1 Quantum simulation on a superconducting quantum processor
Quantum simulations enable the study of the quantum dynamical evolution of the Hubbard model. Using the model in Fig. 4(a) as an example, we illustrate how such research can be conducted on a superconducting quantum chip. A 14-qubit chip can be constructed, as shown in Fig. 4(c), consisting of two 6-qubit rings forming inner and outer circles, with two additional qubits connected to the outer ring. Between neighboring qubits within each ring, there is a nearest-neighbor XY interaction, while between corresponding qubits in the inner and outer rings, there is a ZZ interaction. The other two qubits have only ZZ coupling with the qubits in the outer ring. The $XY$ interaction term corresponds to the hopping between neighboring lattice sites in our model, while the $ZZ$ interaction represents the repulsive or attractive interaction between different particles at the same site. Interaction parameters between qubits on the superconducting qubit chip can be tuned via couplers. Additionally, the coupling strength and sign between superconducting qubits can be controlled through these couplers, enabling quantum simulation of the $e^{i\hat{H}t_{i}}\hat{O_{i}}e^{-i\hat{H}t_{i}}$ sequence evolution. There have been some experimental demonstrations of dynamical evolution in one-dimensional bosonic Hubbard models on superconducting qubit chips [35], as well as studies on thermalization and scrambling in bosonic Hubbard models on ladder structures [24, 23], and other experiments on two-dimensional bosonic Hubbard models [36, 37, 38].